A current of 5.00 Amperes is maintained in a simple circuit with a total resistance of 3.50 ohms. How much heat is generated in 2.50 s?
Answer #1

P = I2R = 52 * 3.5Ω = 87.5 w = 87.5 j/s

87.5j/s * 2.5s = 218.8J